OpenStudy (anonymous):

Derivative of y= (1+x^1/2)^2

OpenStudy (michele_laino):

here you have to apply the formula of the derivative of a function of function, namely: \[\frac{{df}}{{dx}} = 2\left( {1 + {x^2}} \right) \cdot \frac{{d\left( {1 + {x^2}} \right)}}{{dx}}\]

OpenStudy (sweetburger):

chain rule right? take the exponent bring it the front then leave that and then proceed to take the derivative of whatever is inside of the parenthesis

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

oops.. I made an error, here is the right step: \[\frac{{dy}}{{dx}} = 2\left( {1 + \sqrt x } \right) \cdot \frac{{d\left( {1 + \sqrt x } \right)}}{{dx}}\]

OpenStudy (anonymous):

Thank you! I understand it now